the pressure of a gas at constant temperature is tripled, if the final volume was 120cm3. find the initial volume of the gas
1
Expert's answer
2020-04-22T09:19:36-0400

At constant temperature: P1V1 = P2V2

V1 = P2V2/P1

P2 = 3P1

So

V1 = 3P1V2/P1 = 3V2 = 3*120 = 360 (cm3)